Solve by substitution:
x+3y=-25
y=-5x

Answer by Prithwis(166)

x+3y=-25
y=-5x
x+3(-5x)=-25
x-15x=-25
-14x=-25
x=25/14
y=-5(25/14)
y=-125/14
(x,y) = (25/14,-125/14)